OpenStudy (anonymous):

use implicit differentiation to find dy/dx. 2xy-y^2=1. answers are: x/(y-x), x/(x-y), y/(y-x), or y/(x-y).

hartnn (hartnn):

Product rule for 2xy Chain rule for y^2 know those rules already ?

OpenStudy (anonymous):

yes, but could you show your steps, please? They're new to me...

hartnn (hartnn):

let me give you an example instead, say if i have \(x^2*e^y\) using product rule, its derivative will be \((x^2)' e^y + x^2 (e^y)' \\ = 2x e^y + x^2 e^y \dfrac{dy}{dx}\) see if you get this. i have also used chain rule for 2nd term.

hartnn (hartnn):

if you get that, then derivative of xy will be as easy as eating an icecream :P :)

OpenStudy (anonymous):

so i would 2x(dy/dx)+y(2)-2y=0?

hartnn (hartnn):

"2x(dy/dx)+y(2)" is exactly correct for the derivative of 2xy! for the derivative of y^2, we will need chain rule. like derivative of y^8 = 8y^7 dy/dx

OpenStudy (anonymous):

oh, i left off the dy/dx! so the derivative for y^2 would be 2y(dy/dx)?

hartnn (hartnn):

yes! :) 2x(dy/dx)+y(2)-2y(dy/dx)=0 now just isolate dy/dx

OpenStudy (anonymous):

so subtract 2y to the left and factor dy/dx. (dy/dx)(2x-2y)=-2y then divide both sides by 2x-2y which equals -y/x-y. then i multiply by -1/-1 and i get y/y-x. right?

hartnn (hartnn):

it surely is correct :) good work!

OpenStudy (anonymous):

Thank you so much!

hartnn (hartnn):

most welcome ^_^